OpenStudy (anonymous):

PLEASE HELP!!!! I CANNOT FIGURE THIS OUT FOR THE LIFE OF ME!!! After isolating the sulfonated product, the filtrate volume is found to be 19.47 mL. A dilution using 1 mL of filtrate and 99 mL of water gives a solution with an absorbance of 0.743. How many moles of indigo carmine were produced? (Molar absorptivity of indigo carmine was found to be 12,100 M-1cm-1)

OpenStudy (anonymous):

??????

OpenStudy (aaronq):

Beer-Lambert Law: \(\sf Abs=\epsilon*l*M\) \(\sf \epsilon\)=molar absorptivity coefficient \(l\)= path length \(M\)= Molarity (concentration) ; \(M=\dfrac{n_{solute}}{L_{solution}}\) (where n=moles) substitute in: \(\sf Abs=\epsilon*l*\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

I realize this is the right equation to use, but do I use the 100mL or the 25.65mL?

OpenStudy (aaronq):

first find the concentration in the diluted sample, 1 mL filtrate + 99 mL water =100 mL once you know the concentration scale that up to the volume produced. (dont use the extended form of the formula i wrote)

OpenStudy (anonymous):

also, do we need to do anything with the cm^-1?

OpenStudy (aaronq):

nope, that accounts for cuvette, most cuvettes are 1 cm in length. \(l=1~cm\)

OpenStudy (anonymous):

right, I forgot about that. So I put Moles=(.616)(.1L)/(12100M^-1cm^-1) but I still got it wrong.

OpenStudy (aaronq):

where did you get 0.616?

OpenStudy (anonymous):

sorry, I re-opened the problem, and it gave me a different number than .743

OpenStudy (aaronq):

oh true. well by doing that you're only finding the concentration of the diluted sample. you have to find the concentration of the original sample, use: \(M_1V_1=M_2V_2\) Use the molarity and the volume to find the moles.

OpenStudy (anonymous):

I was thinking that, that we were using the volume of the diluted solution, but didn't think it mattered... THANKS!!!

OpenStudy (aaronq):

no problem!

OpenStudy (anonymous):

So just to check, I would use the answer from beers law, and then solve for M1=(ANS)(.1L)/(19.47/1000) using the numbers from this problem

OpenStudy (aaronq):

yeah, then to get moles multiply the volume and molarity